OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y - 2)2 = 9

OpenStudy (anonymous):

Equation of a circle has its center at (h, k) and a radius of a units is \[ (x - h)^2 + (y - k)^2 = a^2 \]

OpenStudy (anonymous):

Hence, \[ (x + 1)^2 + (y - 2)^2 = 9 \]

OpenStudy (anonymous):

Try to think of the definition of a circle: The set of all points that are equidistant from a given point.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Glad to help :)